niahmk:

)Lisa has to cut her grandma's grass this weekend and wants to know exactly how much area she will be cutting. Calculate the area of the polygon. Be sure to show all your work and explain your answer. Six-sided polygon that includes two isosceles right triangles, one with height and base of 20 feet, the other height and base of 6 feet, and one rectangle measuring 30 feet by 6 feet.

new2luv:

do you know how to find the area of a polygon

niahmk:

30 × 6 i think

new2luv:

so what is that

new2luv:

@niahmk wrote:
30 × 6 i think
and i think your right

niahmk:

@new2luv wrote:
so what is that
180

new2luv:

@niahmk wrote:
@new2luv wrote:
so what is that
180
yes

Florisalreadytaken:

what what what? Six-sided polygon that includes `two isosceles right triangles`, one with height and base of 20 feet , the other height and base of 6 feet, `and one rectangle` measuring 30 feet by 6 feet. i dont think an image would be necessary for this we have to find the area of the shapes we're given since we're given the triangles first, we will start off with that \[ \color{dodgerblue}{\triangle}\color{lightskyblue}{A} \color{lightskyblue}{\scriptsize\text{rea}} =\frac{\color{chocolate}{b} \times \color{teal}{h}} {2} \] for the 1st triangle we have \( \color{chocolate}{b} = 20 \) and \( \color{teal}{h} = 20 \) for the 2nd triangle we have \( \color{chocolate}{b} = 6 \) and \( \color{teal}{h} = 6 \)

Florisalreadytaken:

after you do that, fid the area of the rectangle \[ \color{dodgerblue}{ \sqsubset \! \sqsupset }\color{lightskyblue}{A} \color{lightskyblue}{\scriptsize\text{rea}} = \color{chocolate}{b} \times \color{teal}{h} ​\] we're given that \( \color{chocolate}{b} = 30 \) and \( \color{teal}{h} = 6 \) -- it might also be vice versa, but it doesnt really matter after you fo the math, sum it all up together, and there is your answer